OpenStudy (anonymous):

Two numbers are in the ratio 5:2. One half of their sum is 10.5. Find the numbers.

OpenStudy (anonymous):

well call the numbers x and y. you know 2x = 5y, right?

OpenStudy (anonymous):

and what second equation can you write that describes their sum?

OpenStudy (anonymous):

I am not sure about the second equation

OpenStudy (anonymous):

what if I told you that the sum of the two numbers was 10.5, could you write that as an equation?

OpenStudy (anonymous):

x+y=10.5

OpenStudy (anonymous):

yes! so now, instead of "the sum of x and y", how can you write "half the sum of x and y"?

OpenStudy (anonymous):

.5x+.5y=10.5

OpenStudy (anonymous):

yep, that's it. now just solve that simultaneously with the first equation, 2x = 5y

OpenStudy (anonymous):

I don't know how can u show me I haven't learned how to solve 2 variable equations

OpenStudy (anonymous):

well if you multiply that whole equation through by 2 you get 5x + 5y = 21 then isolate x 5x = 21-5y x = (21-5y)/5 Now you know x in terms of y, so substitute that expression (21-5y)/5 into the other equation 2x = 5y in place of x. You'll end up with a single equation with a single variable.

OpenStudy (anonymous):

oh that make sense what do you do after

OpenStudy (anonymous):

solve for x, then plug that value into either equation to find y

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

it does not give a whole number

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

but the answer in the book is 15 & 6 but how

OpenStudy (anonymous):

hi i'm back

OpenStudy (anonymous):

so you know: 2x = 5y 0.5x + 0.5y = 10.5 multiply second equation by 2, you get x + y = 21 isolate y, you get y = 21-x now substitute that into the second equation: 2x = 5y 2x = 5(21-x) 2x = 105 - 5x 7x = 105 x = 15 now use the original equation again: x + y = 21 15 + y = 21 y = 6